# Two statistical problems for multivariate mixture distributions

### v1

@@ 2.4. Strong sm-uniqueness sets
3. Comparing random partitions by model-based methods
3.1. Introduction
@@
The problem of comparing two different partitions of a finite set of observations is an important classical problem in the clustering literature. There is a large amount of work on the subject with several different proposals.
@@
Instead of measuring agreement between two different partitions, we consider the problem of measuring the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures), when both distributions correspond to finite mixtures of multivariate normal or $t$-distributions. In particular, we are able to treat high-dimensional data, and to perform a bootstrap test for $H0$ being that both partitions have been generated by the same mixture distribution against $HA$ that the partitions have been generated by two different mixture distributions.
@@
3.2. Description of the algorithms
3.2.1. Two independent samples
@@
Suppose that we want to compare the results obtained from two clustering procedures but from the same variables. We have two independent samples $\aleph_{1}=\{X_{1},\dots X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{m}\}\subset\mathbb{R}^{d}$, where $\aleph_{1}$ and $\aleph_{2}$ are iid samples with distributions $P,Q$ respectively, both being mixtures of multivariate normal or $t$-distributions.
We proceed as follows. Let $P_{\ell}$ and $Q_{m}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Then we apply Theorem 2.1 (or Theorem 2.2 respectively), and report as a measure of discrepancy between the clusters the average of the projected Kolmogorov–Smirnov distance between the empirical distributions $P_{\ell}$ and $Q_{m}$, along the set of random directions given in Theorem 2.3.
More precisely, we perform the following algorithm.
Algorithm 3.1.
Fix an integer $k\geq\frac{1}{2}(l+m-1)(d^{2}+d-2)-1$.
(i) Take $k$ independent random directions $u_{1},\dots,u_{k}$, uniformly distributed in the unit sphere of $\mathbb{R}^{d}$. According to Theorem 2.3, the set $\{u_{1},\ldots,u_{k}\}$ is a strong sm-uniqueness set for $\mathbb{R}^{d}$ with probability one.
(ii) For each direction $u_{i}$, project the data $\aleph_{1}$ and $\aleph_{2}$ onto the direction $u_{i}$, and thereby obtain two univariate samples
$\mathcal{M}_{1i}$ =$=\{Z_{1i}=\langle u_{i},X_{1}\rangle,\dots,Z_{\ell i}=\langle u_{i},X_{\ell}\rangle\},$
$\mathcal{M}_{2i}$ =$=\{W_{1i}=\langle u_{i},Y_{1}\rangle,\ldots,W_{mi}=\langle u_{i},Y_{m}\rangle\}.$
(iv) For each $i\in\{1,\dots k\}$, calculate the Kolmogorov–Smirnov distance between the empirical distributions associated with the samples $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$. Denote it by $\textit{KS}(i)$.
(v) Output: report as a measure of agreement $\textit{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\textit{KS}(i)$.
Remark.
Algorithm 3.1 can also be used for the two-sample testing problem for mixtures of Gaussian or $t$-distributions. Indeed, given the two samples $\aleph_{1}$ and $\aleph_{2}$ we consider the statistical test problem: $H0:P=Q$ vs $HA:P\neq Q$. Next we perform the algorithm and calculated $\textit{MA}_{k}$. Once $\textit{MA}_{k}$ has been calculated, the critical value can be found using bootstrap as in [11].
3.2.2. One-sample and different clustering procedures
Suppose now that we are given one sample =$\aleph=\{X_{1},\dots,X_{N}\}\subset\mathbb{R}^{d}$. We split the sample into two disjoint subsets $\aleph_{1},\aleph_{2}$. We apply to each subsample a different clustering procedure. For instance, we apply GMM with different prior distributions.
A first naive approach to the problem could be as follows. For each of the the clusters outputs, we want to define a distribution that corresponds to a Gaussian-mixture model. We build it up by considering the mean and the covariance matrix of each cluster, and the weights given by the proportion of points in each cluster. Let $\widehat{P}_{1},\widehat{P}_{2}$ denote those distributions. Robust estimators of the means and covariance matrices can be considered also. Next we generate two samples of the same size, one with distribution $\widehat{P}_{1}$ and the other one with distribution $\widehat{P}_{2}$. Finally we apply Algorithm 3.1 to these samples.
However this approach is too complicated, in particular for high-dimensional data. Since we will end up working with one-dimensional projections, we propose instead the following simpler algorithm. First we observe that, if the data is projected using a linear transformation and a GMM is fitted on the original data, then the density on the projection subspace is again a mixture of Gaussian variables with covariance matrices and mean vectors obtained by a linear transformation of the parameters estimated in the original scale of the data. Thus, for each random direction, we can fit a univariate Gaussian mixture of the projected data. We estimate for each direction the mean and the variance of the projected data, and we use the weights given by the proportion of points at each cluster.
More precisely, we perform the following algorithm:
Algorithm 3.2.
We start with two partitions outputs of the clustering procedures: $\{C_{1},\dots,C_{\ell}\}$ and $\{E_{1},\ldots,E_{m}\}$.
(i) Let $\alpha_{1},\ldots,\alpha_{\ell}$ the proportion of points in the sample at each cluster with procedure $1$, and $\beta_{1},\ldots,\beta_{m}$ the same for procedure $2$.
(ii) For each random direction $u_{i}\leavevmode\nobreak\ (i=1\dots k)$, we consider the projected samples of $C_{j}\leavevmode\nobreak\ (j=1,\dots\ell)$ on the direction $u_{i}$, and obtain estimators of their mean and variance, $\widehat{\mu}_{1ij}$ and $\widehat{\sigma^{2}}_{1ij}$.
(iii) For each random direction $u_{i}\leavevmode\nobreak\ (i=1,\ldots k)$, we consider the projected samples of $E_{j}\leavevmode\nobreak\ (j=1,\dots m)$ on the direction $u_{i}$, and obtain estimators of their mean and variance $\widehat{\mu}_{2ij}$ and $\widehat{\sigma^{2}}_{2ij}$.
(iv) Now consider, for each direction $u_{i}$, the one-dimensional Gaussian mixtures given by
$\widehat{F}_{1i}:=\sum_{j=1}^{\ell}\alpha_{j}N(\widehat{\mu}_{1ij},\widehat{\sigma}^{2}_{1ij})\quad\text{and}\quad\widehat{F}_{2i}:=\sum_{j=1}^{m}\beta_{j}N(\widehat{\mu}_{2ij},\widehat{\sigma}^{2}_{2ij}).$
Using Monte-Carlo, calculate $\textit{KS}(i):=\textit{KS}(\widehat{F}_{1i},\widehat{F}_{2i})$.
(v) Output: report as a measure of agreement $\textit{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\textit{KS}(i)$.
The measure $\textit{MA}_{k}$ takes values between 0 and 1. If the distributions
${F}_{1i}:=\sum_{j=1}^{\ell}\alpha_{j}N(\mu_{1ij},\sigma^{2}_{1ij})\quad\text{and}\quad F_{2i}:=\sum_{j=1}^{m}\beta_{j}N(\mu_{2ij},\sigma^{2}_{2ij}),$
are equal, then $\textit{MA}_{k}=0$, and varies according to the difference between the distribution of the mixtures fitted by both partitions, increasing as long as they are further away.
3.3. Simulations
3.3.1. A simple example of distances between mixtures

### v3

@@ 2.3 Strong sm-uniqueness sets
3 A new estimator for Gaussian mixtures
3.1 Introduction
There is a large literature on this subject. See for instance the well-known work in [54], the recent work [27] and the references therein. The article [54] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}-f_{j}|\geq\epsilon$. In [27], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before [34, Theorem 4.1] below.
3.2 Description of the algorithm
We propose a two-step sequential estimation procedure based on random projections and the empirical characteristic function (ECF).
Let $P$ be a Gaussian mixture on $\mathbb{R}^{d}$, with density
$\sum_{j=1}^{n}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{1/2}}\exp\Bigl{(}-\tfrac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr{)},\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{n}\in[0,1]$, $\sum_{j=1}^{n}\lambda_{j}=1$, $\mu_{1},\dots,\mu_{n}\in\mathbb{R}^{d}$, and $\Sigma_{1},\dots,\Sigma_{n}$ are $d\times d$ positive-definite matrices. We denote $\Lambda:=(\lambda_{1},\dots,\lambda_{n})$, $\bm{\mu}:=(\mu_{1},\dots,\mu_{n})$, $\bm{\Sigma}:=(\Sigma_{1},\dots,\Sigma_{n})$, and $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$.
Given a unit vector $u\in\mathbb{R}^{d}$ and $\theta\in\Theta$, $P_{\langle u\rangle}(\theta)$ denotes the projection of the Gaussian mixture with parameters $\theta$ in the direction $u$. Given an iid sample $X_{1},\dots,X_{N}$, let $P_{N,\langle u\rangle}$ be the projection of the empirical measure in the same direction.
In the first step, the mixture weights $\Lambda$ are estimated by fitting univariate projected mixtures along several random directions. This exploits the fact that linear projections of a Gaussian mixture are again Gaussian mixtures, with transformed means and variances. The estimation is performed by minimizing the discrepancy between the empirical and theoretical characteristic functions of the projected data. Averaging across multiple random directions yields a global estimator of the weights.
In the second step, with the weights fixed, the projected means and variances are re–estimated. From these, the original mean vectors and covariance matrices of the mixture are reconstructed by solving quadratic optimization problems.
The overall procedure is summarized in Algorithm 1.
Algorithm 1 Two-step estimation via random projections and ECF
1:Sample $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
2:Step 1: For each $u_{r}$, estimate $(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)})$ by minimizing $(Z_{u}-F_{u})^{\top}W(Z_{u}-F_{u})$; set $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r}\widehat{\Lambda}_{u_{r},(1)}$
[…]
4:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve equations
$\widehat{\mu}_{j}:=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr{)}^{2},$ (4)
$\widehat{\Sigma}_{j}:=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2},$ (5)
for $j=1,\dots,n$, with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.
[…]
Further details are provided in Appendix A.
Remarks.
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr{|},\ \ \widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr{|},$
for all $j=1,2,\dots,n$.
(iii) The $\operatorname*{argmin}$ in (4) can be easily solved using least squares. To solve the constrained optimization problem of estimating covariance matrices under positive semi-definiteness, see (5), we apply a primal-dual interior-point method (IPM). This approach is well-suited for semidefinite programming (SDP), as it ensures convergence to the global optimum within the cone of symmetric positive semidefinite matrices. The method combines Newtonian steps with a logarithmic barrier function to keep iterates in the feasible region, which in our case corresponds to the set $\mathbb{S}^{+,d}$. This strategy is particularly appropriate given the convexity of our loss function and the structure of the constraint. For background and theoretical guarantees, see [81, 57, 68].
Since $f(\Sigma)=\Bigl{(}\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2}$ is convex and continuous on $\mathbb{S}^{+,d}$, the optimization problem admits a unique solution. Consequently, the penalized objective function used in the IPM Algorithm, $f_{t}(\Sigma)=tf(\Sigma)-\log\det(\Sigma)$, attains a unique minimum over $\mathbb{S}^{+,d}$; see, e.g., [7, 10].
(iv) While semidefinite programming (SDP) is computationally efficient for small-to-moderate dimensions, various alternatives are available for high-dimensional settings. Examples include first-order methods such as the Alternating Direction Method of Multipliers (ADMM), low-rank matrix approximations when the covariance structure allows, and the exploitation of sparsity or block structure in the projection matrices to reduce computational cost, see for example [47] and [79].
(v) In general, the problem becomes increasingly challenging as the dimension of the space grows. The parameter space expands substantially when allowing for general covariance matrices, which increases complexity and renders the algorithm more unstable, since one must optimize over symmetric positive definite matrices in addition to the weights and means. Recently, in [83], a novel and very general problem is studied that includes the more complex setup of compound probability distributions under the restriction of identical and isotropic covariances, which for our setting requires that all covariance matrices be the identity.
(vi) Lastly, we note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
3.3 A consistency result
The results in [76] and [82] that we will use below are based on a general theorem in [43], where the only assumption is that the characteristic function is differentiable. More precisely, it is assumed that the regular case holds, namely that $I_{n}(t)$ can be differentiated under the integral sign. Proposition 1 in [76], which is based on the work in [31], [32], can be stated as follows.
Proposition 8.
Let $\{t_{1},\ldots,t_{m}\}$ be distinct fixed grid points. Then the estimator $\widehat{\theta}$ of $\theta$ is strongly consistent and asymptotically normally distributed, with a given covariance matrix.
How to choose the sequence $\{t_{1},\ldots,t_{m}\}$ as well as the size $m$ in an optimal way for a given sample size $n$ is still an open problem. However, in [32] it is suggested that the sequence should be taken equally spaced, namely $t_{j}=\tau j,\ j=1,\ldots m$ for $\tau\in\mathbb{R}^{+}$.
Theorem 9.
Under the weak assumption that,
$I^{\prime}_{n}(\theta)=\int\frac{\partial}{\partial\theta}|\widehat{\psi}_{n}(t)-\psi(t,\theta)|^{2}dP(t),$
Proof.
First observe that, from the results in [76] and [82] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [76, Proposition 1], where strong consistency and asymptotic normality are established.
for the true values of $\mu_{j}$ and $\Sigma_{j}~(j=1,\dots n)$, we have
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr{)}^{2}-\bigl{(}\langle u_{r},\mu_{j}\rangle-\upsilon_{u_{r},j}\bigr{)}^{2}\Bigr{)}\to 0\quad\text{a.s.}$
and
$\widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl{(}\bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr{)}^{2}-\bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\sigma^{2}_{u_{r},j}\bigr{)}^{2}\Bigr{)}\to 0\quad\text{a.s.}$
for all $j=1,\dots,n$. This concludes the proof. ∎
The only assumption in our Theorem 9 is that $I_{n}(t)$ can be differentiated under the integral sign. As mentioned in [43, p. 258], the only random variables excluded are those with characteristic functions vanishing outside an interval depending on the parameter $\theta$ and those for which $\psi$ is not differentiable with respect to $\theta$. These conditions are fulfilled for the univariate mixtures that we have under consideration. Indeed, the differentiation condition follows for instance from [63].
Therefore our strong consistency result is universal, i.e., it holds without further assumptions for Gaussian or $t$-mixtures. Moreover, we only need $(1/2)(2k-1)(d^{2}+d-2)$ random directions, $k$ being the number of elements in the mixture and $d$ the dimension of the space, i.e., linear in the number of components and quadratic in the dimension. No restriction is placed on the sample size.
An interesting related but different approach is given by Moitra and Valiant in [55] for the estimation of the parameters of mixture of multivariate Gaussians (GMM), where they provide an algorithm that has a polynomial running time. To state their main result, we provide some definitions. In what follows, we write $k]:=\{1,2,\dots,k\}$.
Definition 1.
(i) Given two $n$-dimensional GMMs of $k$ Gaussians, $F=\sum_{i}w_{i}N(\mu_{i},\Sigma_{i})$ and $\widehat{F}=\sum_{i}\widehat{w}_{i}N(\widehat{\mu}_{i},\widehat{\Sigma}_{i})$, we call $\widehat{F}$ an $\epsilon$-close estimate for $F$ if there is a permutation function $\pi:[k]\to[k]$ such that, for all $i\in[k]$,
$\|w_{i}-\widehat{w}_{\pi[i]}\|\leq\epsilon\quad\text{and}\quad D(N(\mu_{i},\Sigma_{i}),N(\widehat{\mu}_{\pi(i)},\widehat{\Sigma}_{\pi(i)}))\leq\epsilon.$
(ii) We call a GMM $F=\sum_{i}w_{i}F_{i}$ $\epsilon$-statistically learnable if $\min_{i}w_{i}\geq\epsilon$ and $\min_{i}D(F_{i},F_{j})\geq\epsilon$.
The result of Moitra and Valiant is given in the following theorem.
Theorem 10 ([55, Theorem 1]).
Given an $n$-dimensional mixture of $k$ Gaussians $F$ that is $\epsilon$-statistically learnable, there is an algorithm that, with probability at least $1-\delta$, outputs an $\epsilon$-close estimate $\widehat{F}$, and the running-time and data requirements of the algorithm (for any fixed $k$) are polynomial in $n$, $1/\epsilon$ and $1/\delta$.
If one is only interested on D(F,F^)D($\widehat{F}$), then in [55, Corollary 2] there is stated a similar result without the restriction that $F$ be $\epsilon$-statistically learnable.
Sharp results for the estimation of univariate finite mixtures are also provided by Heinrich and Kahn in [44], where they study the rates of convergence of the parameters of the mixture, and, under some regularity and strong identifiability conditions, find the optimal local minimax rate of estimation of a univariate mixture with $m$ components. However they do not address the multivariate case.
3.4 Simulations
3.4.1 Example 1.
$F$ $=\lambda_{1}\,t(\eta,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\,t(\eta,\mu_{2},\Sigma_{2}),$
$\mu_{1}$ $=(1,0),\quad\mu_{2}=(5,0),$
$\Sigma_{1}$ $=\begin{pmatrix}1&0\\ 0&1/2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}1/2&0\\ 0&1\end{pmatrix}$,
[…]
Fig. 1: Depiction of $1000$ simulations of the mixture $F$.
In this Example, we present a small comparative study of the estimation derived by the random projections method (RP) with respect to the expectation-maximization method for mixtures of t-distributions (we denote EM-st). This last method, developed in [62], is implemented by means of the function fmmst of the EMMIXuskew package of the R language. In this case the skew parameter was prefixed to 0.
Table 1: Averages values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 1 0 1 1/2 0
2 5 0 1/2 1 0
RP 1 0.99 -0.01 0.99 0.49 0
sd (0.10) (0.10) (0.14) (0.13) (0.07)
2 4.99 0.014 0.56 0.73 -0.01
sd (0.13) (0.12) (0.05) (0.19) (0.06)
EM-st 1 1.22 0.07 0.89 0.62 0.01
sd (0.15) (0.16) (0.14) (0.13) (0.09)
2 4.56 -0.10 0.51 1.15 -0.01
sd (0.10) (0.25) (0.07) (0.15) (0.07)
Fig. 2: Comparison between the EM-st and RP methods using the $L^{2}$-distance between the parameters.
3.4.2 Example 2
$F$ $=\lambda_{1}\,N(\mu_{1},\Sigma_{1})+(1-\lambda_{1}-\gamma)\,N(\mu_{2},\Sigma_{2})+\gamma\,\mathcal{C}a(\mu_{3}),$
$\mu_{1}$ $=(1,0),\quad\mu_{2}=(5,0),\quad\mu_{3}=(3,0),$
$\Sigma_{1}$ $=\begin{pmatrix}4&0\\ 0&2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}2&0\\ 0&4\end{pmatrix}$,
[…]
Fig. 3: Comparison between the Robust EM and RP Boxplots when outliers are added, taking $\mu_{3}=(3,0)$.
3.5 An example with real data
Fig. 4: Hexagonal binning for bivariate data. The plane is tessellated by a regular grid of hexagons (cells). Then the hexagons are plotted using an intensity of color proportional to the number of points falling in each cell. Higher socio-economic level for smaller values on the $x$-axis, better mathematical ability for larger values on the $y$-axis.
$\widehat{\Sigma}_{1}=\begin{pmatrix}0.39&0.15\\ 0.15&0.65\end{pmatrix},\quad\widehat{\Sigma}_{2}=\begin{pmatrix}0.16&0.08\\ 0.08&0.48\end{pmatrix}$.
These values are similar to those obtained in the mixture estimation.
4 Comparing random partitions by model-based methods
4.1 Introduction
@@
The problem of comparing two different partitions of a finite set of observations is an important classical problem in the clustering literature. There is a large amount of work on the subject with several different proposals. Our aim in this section is to propose a new approach to this problem. We consider not only the cluster labels assigned by the clustering algorithm but also the distribution of the resulting groups. Note that we do not address the general problem of model-based clustering, which is a broad topic (see [37] and [9]).
@@
Instead of measuring agreement between two different partitions, we consider the problem of measuring the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures), when both distributions correspond to finite mixtures of multivariate normal or $t$-distributions. In particular, we are able to treat high-dimensional data.
@@
4.2 Description of the algorithm
@@
Suppose that we want to compare the results obtained from two clustering procedures, both applied to the same set of variables. We consider two independent samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$, with distributions $P$ and $Q$ respectively, both being mixtures of $m$ multivariate Gaussian or $t$-distributions.
Our method relies on random projections and Kolmogorov–Smirnov (KS) distances. Specifically, each sample is projected onto several random one-dimensional subspaces, empirical distributions are computed, and the KS distances are evaluated between the projected samples. Aggregating these distances provides a measure of discrepancy between the two clustering results.
We consider two iid samples
$\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}\ \mbox{and}\ \aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d},$
drawn from distributions $P$ and $Q$, each being a mixture of $m$ multivariate Gaussian or $t$-distributions. Let $P_{\ell}$ and $Q_{r}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Applying Theorems 5 and 6, the discrepancy between the clusters is measured by the average of the projected KS-distances between $P_{\ell}$ and $Q_{r}$ along a strong sm-uniqueness set of random directions (Theorem 7).
For each direction $u_{i}$, denote the projected samples by
$\mathcal{M}_{1i}=\{\langle u_{i},X_{j}\rangle\}_{j=1}^{\ell},\qquad\mathcal{M}_{2i}=\{\langle u_{i},Y_{j}\rangle\}_{j=1}^{r},$
with corresponding empirical CDFs $F_{1i}$ and $F_{2i}$. The one-dimensional Kolmogorov–Smirnov statistic is
$\mathrm{KS}(i)=\sup_{t\in\mathbb{R}}|F_{1i}(t)-F_{2i}(t)|.$
We then define
$D_{k}:=\max_{1\leq i\leq k}\mathrm{KS}(i).$
As $\min(\ell,r)\to\infty$, the asymptotic distribution of $D_{k}$ satisfies
$\lim_{\min(\ell,r)\to\infty}\Pr(D_{k}\leq t)=\lim_{\min(\ell,r)\to\infty}\Big{[}\Pr(\mathrm{KS}(1)\leq t)\Big{]}^{k},$
which provides the basis for an asymptotic test. A drawback is the loss of the distribution-free property, as the distribution of $D_{k}$ depends on the covariance structure of the underlying processes.
The overall scheme is summarized in Algorithm 2.
Algorithm 2 Agreement measure via random projections and Kolmogorov–Smirnov distances
1:Integers $m,d$; samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$, $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)-1$
2:Draw $k$ random directions $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
3:for $i=1$ to $k$ do
4:  Project samples: $\mathcal{M}_{1i}=\{\langle u_{i},X_{j}\rangle\}_{j=1}^{\ell}$,  $\mathcal{M}_{2i}=\{\langle u_{i},Y_{j}\rangle\}_{j=1}^{r}$
5:  Compute empirical CDFs $F_{1i},F_{2i}$ from $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$
6:  $\mathrm{KS}(i)\leftarrow\sup_{t\in\mathbb{R}}|F_{1i}(t)-F_{2i}(t)|$
7:end for
8:return $\mathrm{D}_{k}\leftarrow\max_{i=1,\dots,k}\mathrm{KS}(i)$
Remarks.
(i) An alternative measure is the average
$\mathrm{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i),$
where $\mathrm{KS}(i)$ is the KS distance between the empirical distributions of $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$. If $F_{i}$ and $G_{i}$ are the true projected distributions, then by the triangle inequality,
$0\leq\mathrm{KS}(i)\leq\mathrm{KS}(F_{\mathcal{M}_{1i}},F_{i})+\mathrm{KS}(F_{i},G_{i})+\mathrm{KS}(G_{\mathcal{M}_{2i}},G_{i}),$
which implies
$\lim_{\min(\ell,r)\to\infty}\mathrm{MA}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(F_{i},G_{i}).$
Theorems 5 and 6 show that if this limit is zero, then $P=Q$. However, the asymptotic distribution of $\mathrm{MA}_{k}$ is unknown, since projections onto different directions are only conditionally independent.
(ii) If the sample sizes $\ell$ and $r$ are large, one may split them into $k$ disjoint subsamples and use different subsamples for each direction $u_{i}$. In this case, $\mathrm{MA}_{k}$ (or $D_{k}$) becomes an average (or maximum) of $k$ independent one-dimensional KS statistics, which is distribution-free. This strategy, however, comes at the cost of reduced statistical power.
4.3 Simulations